help solve number 12 plz
Question 12 10 pts What is the pH of 0.1 MNH4Cl in solution? The Kb of NH3 is 1.8 x 105. Answer to 2 decimal places please.
Answer #1

For calculating the P, eventually want [H0]*. So determine the Ka of NH3 from Kb. K, K 1x10-14 1.8x10-5 =5.56x10-10 HydrolysiNH4+ + H2O + NH3 + H30+ Calculate [H3O+] using the Ka equation K - NH3][H0]* [NH] _[H,0*7 [NH, ] 5.56 x10-10 [H,0^] - 5.56 [Then, pH = -log [H3O+] pH = -log 7.45*10^-5 pH = 4.12